Calculate the molecular weight of the phenol-formaldehyde (Bakelite) repeatini 73 amu 0 100 amu 88.7 amu 133 amu
Answer #1

Repeating unit of phenol formaldehyde (Bakelite is (64502 Hence; molecular = 1288+ 5X1 +32 = 96+5 +32 2133 amy e molecular wePlease like this answer. THANK YOU

Please comment if you have any doubts about this answer.

All the best for your future.